OpenStudy (anonymous):

Factor completely: 2x2 + 20x + 50

OpenStudy (anonymous):

2. 2x^2 - 20x + 50 = 2(x^2 - 10x + 25) = 2(x-5)^2

OpenStudy (anonymous):

you put - 20x and its suppose to be + so is all of it the same but i just add + ?

OpenStudy (anonymous):

2(x+5)^2